# A Survey on Adversarial Contention Resolution

### v3

@@ 3.3.2. The Wakeup Problem
Under no-CD, Jurdziński and Stachowiak (jurdzinski2005probabilistic, ; jurdzinski2002probabilistic, ) present randomized algorithms with probability of error $\epsilon$ for solving the wakeup problem under different settings: where participants are labeled or unlabeled, where there is a global clock or not, and where $n$ is known or unknown. We note that, while deterministic approaches rely on the use of labels, this is an example of labels being used with a randomization solution; in this case, the labels are used to approximate system size. The authors are concerned with uniform algorithms (recall Section 2.3). Perhaps unsurprisingly, aspects of these wakeup algorithms are reminiscent of exponential backoff. One of the main findings is a near-exponential separation between the global and local clock settings when $n$ is unknown and there no labeling; the former can be solved with latency $O(\log(n)\log(1/\epsilon))$, while the latter requires $\Omega(n/\log n)$ latency. In the cases where $n$ is known or the participants are labeled (or both), the run time is $O(\log n\log(1/\epsilon))$. For uniform algorithms, the authors also show a nearly-tight lower bound of $\Omega\left(\frac{\log(n)\log(1/\epsilon)}{\log\log(n)+\log\log(1/\epsilon)}\right)$ in the case where $n$ is known and the participants are labeled. As stated earlier in Section 2.3, this bound improved by Farach-Colton et al. (farach-colton:lower-bounds, ) to $\Omega(\log(n)\log(1/\epsilon))$, and then later shown to hold for all randomized algorithms by Newport (newport2014radio, ; newport:radio-journal, ).
Under the SINR model (recall Section 2.3), and in the absence of a global clock and assuming participants know a polynomial upper bound on the system size, Jurdziński and Stachowiak (jurdzinski2015cost, ) give a Monte Carlo algorithm that has w.h.p. $O(\log^{2}n/\log\log n)$ latency, and a Las Vegas algorithm with expected $O(\log^{2}n/\log\log n)$ latency. We note that this is better than the deterministic algorithm in (jurdzinski2013distributed, ) of $O(\log^{2}n)$ latency (recall Section 3.2.2).
@@
Summary. For the $k$-selection problem, a key message is that we can get asymptotically optimal $O(k)$ latency. For the wakeup problem, the situation is also significantly improved by randomization. Specifically, compared to the results surveyed in Section 3.2.2, we see that randomization often allows for $O(\texttt{polylog}(n))$ latency in contrast to $\omega(n)$ latency in the single-hop scenario; however, in the local clock setting without knowledge of key parameters, the latency remains high. That said, in comparison to the results from the AQT and deterministic approaches, it is clear that randomization can offer significant improvements.
Another aspect of randomized approaches to the selection problem is the focus on the number of access attempts, which translates to energy consumption. The $O(\log\log^{*}n)$ complexity achieved by Bender et al. (bender:contention, ) exploits the ability to send and receive messages of $\Theta(\log n)$ bits. Under the more standard ternary feedback model, the best results achieve a polylogarithmic in $n$ number of access attempts.
Finally, we note that while many of the randomized results for contention resolution achieve asymptotically-optimal latency, they are fragile in the sense that even a small amount of channel interference can cause the corresponding algorithm to fail. For example, the result of Bender et al. (bender:contention, ) estimates the number of packets contending for channel access; even a single slot with interference can cause a gross overestimate. Another instance is a result by De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ), which depends on coordination messages issued by a leader; such messages can be disrupted by interference. The algorithm by Bender et al. (bender2020contention, ) is also unable to maintain asymptotically-optimal latency in the face of jamming. Thus, robustness to malicious interference presents a challenge, and it is the subject we turn to next.
@@ 4. Jamming-Resistant Contention Resolution
A jamming attack is characterized by deliberate collisions on the multiple access channel. By deviating from a prescribed protocol, a jamming device (jammer) can launch a denial-of-service attack, preventing other devices from being able to communicate (DBLP:reference/algo/RichaS16, ). There have been several well-publicized instances of jamming (FCC:humphreys, ; marriott, ; priest, ; nicholl, ), behavior that may become more common given the trend towards wireless personal devices. For example, alluding to the quote above, Father Madonna installed a store-bought jamming device in his church to prevent attendees from disrupting his services with cellphone usage. However, this device also interfered with nearby wireless devices, with shop owners complaining that credit card payment systems were affected.
There are a number of papers addressing applied security considerations with respect to jamming (for examples, see (bayraktaroglu:performance, ; xu:feasibility, ; deejam:wood, ; aschenbruck:simulative, ; wood:denial, ; xu:jamming, ) and surveys (mpitziopoulos:survey, ; VADLAMANI201676, )). For example, spread spectrum techniques (spread:foiling, ; navda:using, ), rerouting defenses (wood2003jam, ; han2018hierarchical, ), and specialized antenna technology (zheng2018design, ; mcmilin2015gps, ). From a theory perspective, several fundamental distributed computing challenges—such as broadcast (koo2, ; gilbert:near, ; gilbert:making, ; chen:broadcasting, ), leader election (gilbert:malicious, ; gilbert2009malicious, ), gossip (gilbert:interference, ; dolev:gossiping, ), node discovery (meier:speed, ), and simple point-to-point communication (king:conflict, ; DBLP:journals/dc/KingPSY18, )—can be solved on a shared channel despite jamming (see (young:overcoming, ; Bender:2015:RA:2818936.2818949, )). Here, we focus on prior contention-resolution results that offer provable guarantees on throughput or latency, despite jamming.
@@
A jamming adversary poses at least two significant new challenges over the dynamic-arrivals model considered in Section 3. First, it may not be obvious that non-trivial makespan guarantees are possible; for example, if jamming occurs in every slot in perpetuity, then no participants can succeed. Therefore, we must either constrain the jamming power of the adversary (e.g., by assuming a bounded adversary), or alter our measure of throughput to address this case (e.g., such as notions of competitiveness and implicit throughput). Second, the algorithms surveyed so far that rely on channel feedback to adjust contention—for example, by estimating $n$ (such as in (BenderKPY18, ; DBLP:journals/jacm/GreenbergFL87, )) or by exchanging messages between participants—are vulnerable, since jamming allows the adversary to create collisions and prevent communication.
With regard to the first challenge, a common approach is to measure performance over non-jammed slots; for example, this is done by Chang et al. (ChangJP19, ) and Bender et al. (bender:how, ). Specifically, assuming that at least one participant is present in the system, a slot that is not jammed, but where no participant succeeds, is said to be wasted. Guaranteeing that only a small constant fraction of slots are wasted allows us to derive meaningful results despite an adversary that jams aggressively, perhaps even in every slot.
Addressing the second challenge of adjusting the contention requires care, and this is demonstrated by the elegant algorithm by Chang et al. (ChangJP19, ). Participants follow some simple rules for updating the sending probability in each slot: increasing if there is silence, decreasing if there is noise, and leaving it unchanged in the case of a success. The adversary may lower the sending probabilities of participants by jamming (i.e., adding noise), thus depressing contention. However, the update rule allows for contention to quickly recover once the jamming ceases (although, it may resume again later), and few slots are wasted as a result. Notably, this design allows the algorithm to achieve roughly $1/e$ throughput in the non-jammed slots, which rivals that of slotted ALOHA, despite the more-challenging adversarial setting.
@@
An early series of results against a bounded adversary was initiated by Awerbuch et al. (awerbuch:jamming, ; DBLP:journals/talg/AwerbuchRSSZ14, ), who consider the following model. There are $n$ honest participants equipped with CD, and the network is saturated (recall Section 2.1). For some $T\in\mathbb{N}$ and constant $10<\epsilon<$ 1, there is an adversary that can jam at most $(1-\epsilon)w$ of the slots in any consecutive $w\geq T$ slots. This adversary is referred to as $(T,1-\epsilon)$-bounded, and the value of $\epsilon$ is unknown to participants.
@@
Such bounded jamming may be well-motivated in practice, where aggressive jamming can reveal that an attack is occurring, as well as the location of the malicious device(s) (for examples, see (liu2009localizing, ; liu2013error, ; xu:feasibility, )), or correspond to a limited power supply. The adversary is also adaptive, meaning that the adversary knows the full history of the actions performed by honest nodes for all the previous slots, and may use this information to inform its jamming strategy in the current slot; however, the adversary does not know what actions participants are taking in the current slot. When the adversary jams a slot, all listening participants hear jamming; this is referred to as a 1-uniform adversary.
@@
Awerbuch et al. (awerbuch:jamming, ; DBLP:journals/talg/AwerbuchRSSZ14, ) define an algorithm to be constant competitive if, for any sufficiently large number of slots, the honest nodes perform successful message transmissions in a constant fraction of the non-jammed steps. The authors provide a randomized algorithm that is constant competitive with high probability (w.h.p.) in $n$, when all participants know a common parameter, $\gamma$$=O(1/(\log T+\log\log n))$, and the algorithm is run for a sufficiently large number of slots. Roughly speaking, in a slot, the algorithm dictates that each participant $u$ sends a message with probability $p_{u}$. If $u$ does not send, it listens and updates $p_{u}$ via a multiplicative factor which can increase or decrease, depending on whether or not the slot is idle. Notably, this multiplicative weights update method shares common ground with the later algorithm by Chang et al. (ChangJP19, ) discussed above.
@@
Subsequently, this line of work has been extended to several challenging settings, showing constant competitiveness so long as $\gamma$ is known to all the participants, and the algorithm is executed for a sufficiently large number of slots. Richa et al. (richa:jamming2, ) consider a multi-hop network modeled as a unit disk graph (UDG). Informally, under the UDG model, each participant can broadcast within a disk centered about its location, and two participants may communicate if their respective disks intersect (see Schmid and Wattenhofer (schmid2006algorithmic, ) for discussion of the UDG model). First, Richa et al. (richa:jamming2, ; DBLP:journals/dc/RichaSSZ13, ) consider the setting where the adversary is 1-uniform and the UDG is connected, or each participant has at least $2/\epsilon$ neighbors. Then, against a $1$-uniform $(T,1-\epsilon)$-bounded adversary, the authors provide an algorithm, named Jade, that is w.h.p. constant competitive. Second, Richa et al. (richa:jamming2, ) prove that, if the graph is not connected, or the adversary is $2$-uniform—for any slot, the adversary may choose any partition of the $n$ participants and have the members of one partition hear jamming, while the others do not—and participants each have $o(1/\epsilon)$ neighbors, then Jade can fail to be constant competitive. Against a reactive jamming adversary—who can learn whether or not the channel is idle in the current slot and can choose to jam based on this information—Richa et al. (richa:jamming3, ; DBLP:journals/ton/RichaSSZ13, ) show that constant competitiveness is still possible.
@@
A different approach is taken by Bender et al. (bender:how, ; BenderFGY19, ), who address an adversary that is not constrained in its jamming. A packet starts in an “inactive” state by listening to the channel, and if a sufficient number of empty slots are heard, the packet becomes “active” in the sense of making transmission attempts. This “throttling” mechanism gives time for contention to decrease, thus allowing active packets to succeed. An active packet $u$ will make transmission attempts, but in contrast to updating based on feedback in each slot, $u$ monotonically reduces its sending probability as roughly $1/s_{u}$, where $s_{u}$ is the number of slots since $u$ became active (i.e., the age of $u$). Packet $u$ also listens (when not transmitting), and if a sufficient number of empty slots are heard, it resets $s_{u}=1$. Here, they show expected constant throughput, with each packet making an expected number of transmission attempts that is polylogarithmic in $n+J$, where $J$ is the total number of jammed slots during which there exists at least one packet.
Recently, Bender et al. (bender:fully, ) showed how to achieve $\Omega(1)$ implicit throughput (recall Section 1.1.3) with $O({\texttt{poly}({\log n})})$ channel accesses (both sending and listening), despite an adaptive jamming adversary. This result applies in the channel setting with CD (i.e., ternary channel feedback), but with no control bits allowed, unlike the energy-efficient result in (BenderKPY18, ). The algorithm, named Low-Sensing Backoff, is based on multiplicative weights updates, and in that regard, it shares common ground with some prior results (such as (ChangJP19, ; richa:jamming2, ; richa:jamming3, ; richa:jamming4, ; awerbuch:jamming, ; ogierman2018sade, )). However, in comparison to these results, Low-Sensing Backoff greatly limits the frequency at which packets can sense the channel in order to update, while the lack of control bits greatly limits the amount of information they receive in each update. Analogous results are given for the cases where the number of packets is infinite, and where the adversary is reactive.
@@
Jurdziński et al. (DBLP:conf/waoa/JurdzinskiKL14, ) also consider a single sender and receiver, but in contrast to a parameterized adversary, they consider: an unbounded adversary, and packets of different sizes may arrive over time. If the adversary jams at any time during which a packet is transmitted, the transmission fails and the sender learns of this immediately. The authors derive results on the ratio between the throughput achieved by the algorithm and the throughput achieved by the best scheduling under identical arrival and error patterns; this is referred to as relative throughput. The authors give a deterministic online algorithm that achieves optimal relative throughput.
@@
An impossibility result is given by Bender et al. (bender2020contention, ), who show that with no-CD, no algorithm can guarantee a packet success in time linear in the system size with better than constant probability in the presence of jamming; this result holds even if against a non-adaptive adversary. Subsequently, Chen et al. (DBLP:conf/podc/ChenJZ21, ) give an algorithm in the no-CD case that achieves a trade-off between the throughput and the amount of jamming tolerated. In the challenging case where a constant fraction of slots are jammed, $O(1/\log n)$ throughput is shown, which aligns with the lower bound of Bender et al. (bender2020contention, ).
@@
Summary. Perhaps unsurprisingly, randomization along with CD can provide constant throughput (in non-jammed slots) despite an unbounded adversary; even without CD, jamming can be tolerated, although throughput decreases slightly to $\Theta(1/\log n)$. When the adversary is bounded, a range of results exist. Under a saturated model, randomized results demonstrate constant throughput for a variety of network settings against a bounded adversary, so long as some weak information about the adversary’s power and $n$ is known. In contrast, deterministic results in the AQT setting are more nuanced, being parameterized by aspects of jamming rate and burstiness, and generally more pessimistic.
@@
Existing results that tolerate jamming exhibit a sharp trade-off between the amount of listening devices perform and throughput achieved. On one hand, if each participant listens to the channel in every slot (except when sending), then constant throughput (in expectation or w.h.p. in non-jammed slots) is possible (bender:how, ; BenderFGY19, ; ChangJP19, ; awerbuch:jamming, ; richa:jamming2, ; richa:jamming3, ). Note that such results are not energy efficient, given that listening is roughly as costly as sending in practice (recall Section 1.1.3). On the other hand, if devices never listen to the channel, only sub-constant throughput is possible, and the best result is $O(1/\log n)$ throughput w.h.p. (DBLP:conf/podc/ChenJZ21, ). The recent result in (bender:fully, ) demonstrates that jamming-resistant contention resolution is possible with constant throughput and a polylogarithmic number of channel accesses (both sending and listening) without using control bits. However, it is unclear whether better energy-efficiency bounds are possible in the presence of jamming, perhaps under models with additional feedback permitted on the channel.

### v4

@@ 3.3.2. The Wakeup Problem
Under no-CD, Jurdziński and Stachowiak (jurdzinski2005probabilistic, ; jurdzinski2002probabilistic, ) present randomized algorithms with probability of error $\epsilon$ for solving the wakeup problem under different settings: where participants are labeled or unlabeled, where there is a global clock or not, and where $n$ is known or unknown. We note that, while deterministic approaches rely on the use of labels, this is an example of labels being used in a randomized solution; in this case, the labels are used to approximate system size. The authors are concerned with uniform algorithms (recall Section 2.3). Perhaps unsurprisingly, aspects of these wakeup algorithms are reminiscent of exponential backoff. One of the main findings is a near-exponential separation between the global and local clock settings when $n$ is unknown and there no labeling; the former can be solved with latency $O(\log(n)\log(1/\epsilon))$, while the latter requires $\Omega(n/\log n)$ latency. In the cases where $n$ is known, or the participants are labeled (or both), the run time is $O(\log n\log(1/\epsilon))$. For uniform algorithms, the authors also show a nearly-tight lower bound of $\Omega\left(\frac{\log(n)\log(1/\epsilon)}{\log\log(n)+\log\log(1/\epsilon)}\right)$ in the case where $n$ is known and the participants are labeled. As stated earlier in Section 2.3, this bound improved by Farach-Colton et al. (farach-colton:lower-bounds, ) to $\Omega(\log(n)\log(1/\epsilon))$, and then later shown to hold for all randomized algorithms by Newport (newport2014radio, ; newport:radio-journal, ).
Under the SINR model (recall Section 2.3), and in the absence of a global clock and assuming participants know a polynomial upper bound on the system size, Jurdziński and Stachowiak (jurdzinski2015cost, ) give a Monte Carlo algorithm that has w.h.p. $O(\log^{2}n/\log\log n)$ latency, and a Las Vegas algorithm with expected $O(\log^{2}n/\log\log n)$ latency, which may be contrasted with the deterministic algorithm in (jurdzinski2013distributed, ) of $O(\log^{2}n)$ latency (recall Section 3.2.2).
@@
Summary.
For the $k$-selection problem, a key message is that we can get asymptotically optimal $O(k)$ latency using adaptive algorithms. For the wakeup problem, the situation is also significantly improved by randomization. Specifically, compared to the results surveyed in Section 3.2.2, we see that randomization often allows for $O(\texttt{polylog}(n))$ latency in contrast to $\omega(n)$ latency in the single-hop scenario; however, in the local clock setting without knowledge of key parameters, the latency remains high. That said, in comparison to the results from the AQT and deterministic approaches, it is clear that randomization can offer significant improvements.
Another aspect of randomized approaches to the $k$-selection problem is the focus on the number of access attempts, which translates to energy consumption. The $O(\log\log^{*}n)$ complexity achieved by Bender et al. (bender:contention, ) exploits the ability to send and receive messages of $\Theta(\log n)$ bits. Under the more standard ternary feedback model, the best results achieve a polylogarithmic in $n$ number of access attempts.
Finally, we note that while many of the randomized results for $k$-selection achieve asymptotically-optimal latency, they are fragile in the sense that even a small amount of channel interference can cause the corresponding algorithm to fail. For example, the result of Bender et al. (bender:contention, ) estimates the number of packets contending for channel access; even a single slot with interference can cause a gross overestimate. Another instance is a result by De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ), which depends on coordination messages issued by a leader; such messages can be disrupted by interference. The algorithm by Bender et al. (bender2020contention, ) is also unable to maintain asymptotically-optimal latency in the face of jamming. Thus, robustness to malicious interference presents a challenge, and it is the subject we explore next.
@@ 4. Jamming-Resistant Contention Resolution
A jamming attack is characterized by deliberate collisions on the multiple access channel. By deviating from a prescribed protocol, a jamming device (jammer) can launch a denial-of-service attack, preventing other devices from being able to communicate (DBLP:reference/algo/RichaS16, ). There have been several well-publicized instances of jamming (FCC:humphreys, ; marriott, ; priest, ; nicholl, ), behavior that may become more common given the trend towards wireless personal devices. For example, in reference to the quote above, Father Madonna installed a store-bought jamming device in his church to prevent cellphone usage by attendees from disrupting his services. However, this device also interfered with nearby wireless devices, with shop owners complaining that credit card payment systems were affected.
There are a number of papers addressing applied security considerations with respect to jamming (e.g., (bayraktaroglu:performance, ; xu:feasibility, ; aschenbruck:simulative, ; mpitziopoulos:survey, ; VADLAMANI201676, )); for example, spread spectrum techniques (spread:foiling, ; navda:using, ), rerouting defenses (wood2003jam, ; han2018hierarchical, ), and specialized antenna technology (zheng2018design, ; mcmilin2015gps, ). From a theory perspective, several fundamental distributed computing challenges—such as broadcast (koo2, ; gilbert:near, ; gilbert:making, ; chen:broadcasting, ), leader election (gilbert:malicious, ; gilbert2009malicious, ), gossip (gilbert:interference, ; dolev:gossiping, ), node discovery (meier:speed, ), and simple point-to-point communication (king:conflict, ; DBLP:journals/dc/KingPSY18, )—can be solved on a shared channel despite jamming (see (young:overcoming, ; Bender:2015:RA:2818936.2818949, )). Here, we focus on prior contention-resolution results that offer provable performance guarantees.
@@
A jamming adversary poses at least two significant new challenges over the dynamic-arrivals model considered in Section 3. First, it may not be obvious that non-trivial latency guarantees are possible; for example, if jamming occurs in every slot in perpetuity, then no participants can succeed. Therefore, we must either constrain the jamming power of the adversary (e.g., by assuming a bounded adversary), or alter our measure of throughput to address this case (e.g., via notions of competitiveness and implicit throughput). Second, the algorithms surveyed so far that rely on channel feedback to adjust contention—for example, by estimating $n$ (such as in (BenderKPY18, ; DBLP:journals/jacm/GreenbergFL87, )) or by exchanging messages between participants (such as in (BenderKPY18, ; DeMarco:2017:ASC:3087801.3087831, )—are vulnerable, since jamming allows the adversary to cause overestimates of $n$ and prevent communication.
To address the first challenge, a common approach is to optimize performance in non-jammed slots, as done by Chang et al.(ChangJP19, ) and Bender et al.(bender:how, ). Assuming that at least one participant is present in the system, a slot that is not jammed, but where no participant succeeds, is said to be wasted. Guaranteeing that only a small constant fraction of slots are wasted allows for meaningful results despite an adversary that jams aggressively.
Addressing the second challenge of adjusting the contention requires care, and this is demonstrated by the elegant algorithm of Chang et al. (ChangJP19, ). Participants follow some simple rules for updating the sending probability in each slot: increasing if there is silence, decreasing if there is noise, and leaving it unchanged in the case of a success. The adversary may lower the sending probabilities of participants by jamming (i.e., adding noise), thus depressing contention. However, the update rule allows for contention to quickly recover once the jamming ceases (although, it may resume again later), and few slots are wasted as a result. Notably, this design allows the algorithm to achieve roughly $1/e$ throughput in the non-jammed slots, which rivals that of slotted ALOHA, despite the more-challenging adversarial setting.
@@
An early series of results against a bounded adversary was initiated by Awerbuch et al. (awerbuch:jamming, ; DBLP:journals/talg/AwerbuchRSSZ14, ), who consider the following model. There are $n$ participants equipped with CD, and the network is saturated (recall Section 2.1). For some $T\in\mathbb{N}$ and constant $10<\epsilon<$ 1, there is an adversary that can jam at most $(1-\epsilon)w$ of the slots in any consecutive $w\geq T$ slots. This adversary is referred to as $(T,1-\epsilon)$-bounded, and the value of $\epsilon$ is unknown to participants.
@@
Such bounded jamming may be well-motivated in practice, where aggressive jamming can reveal that an attack is occurring, as well as the location of the malicious device(s) (for examples, see (liu2009localizing, ; liu2013error, ; xu:feasibility, )), or correspond to a limited power supply. The adversary is also adaptive, meaning that the adversary knows the full history of the actions performed by participants for all the previous slots, and may use this information to inform its jamming strategy in the current slot; however, the adversary does not know what actions participants are taking in the current slot. When the adversary jams a slot, all listening participants hear jamming; this is referred to as a 1-uniform adversary.
@@
Awerbuch et al. (awerbuch:jamming, ; DBLP:journals/talg/AwerbuchRSSZ14, ) define an algorithm to be constant competitive if, for any sufficiently large number of slots, participants succeed in a constant fraction of the non-jammed steps. The authors provide a randomized algorithm that is w.h.p. constant competitive when all participants know a common parameter, $\gamma$$=O(1/(\log T+\log\log n))$, and the algorithm is run for a sufficiently large number of slots. Roughly speaking, in a slot, the algorithm dictates that each participant $u$ sends a message with probability $p_{u}$. If $u$ does not send, it listens and updates $p_{u}$ via a multiplicative factor which can increase or decrease, depending on whether or not the slot is idle. Notably, this algorithm design shares common ground with the later work by Chang et al. (ChangJP19, ) discussed above.
@@
Subsequently, this line of work has been extended to several challenging settings, showing constant competitiveness so long as $\gamma$ is known to all the participants, and the algorithm is executed for a sufficiently large number of slots. Richa et al. (richa:jamming2, ) consider a multi-hop network modeled as a unit disk graph (UDG). Informally, under the UDG model, each participant can broadcast within a disk centered about its location, and two participants may communicate if their respective disks intersect (see Schmid and Wattenhofer (schmid2006algorithmic, ) for discussion of the UDG model). First, Richa et al. (richa:jamming2, ; DBLP:journals/dc/RichaSSZ13, ) consider the setting where the $(T,1-\epsilon)$-bounded adversary is 1-uniform, and the UDG is connected or each participant has at least $2/\epsilon$ neighbors. The authors provide an algorithm, named Jade, that is w.h.p. constant competitive. Second, Richa et al. (richa:jamming2, ) prove that, if the graph is not connected, or the adversary is $2$-uniform—for any slot, the adversary selects any partition of the $n$ participants and has the members of one partition hear jamming, while the others do not—and some participants have $o(1/\epsilon)$ neighbors, then Jade can fail to be constant competitive. Against a reactive jamming adversary—who first learns whether the channel is idle in the current slot and then chooses whether to jam—Richa et al. (richa:jamming3, ; DBLP:journals/ton/RichaSSZ13, ) show that constant competitiveness is possible.
@@
A different approach is taken by Bender et al. (bender:how, ; BenderFGY19, ), who address an adversary that is not constrained in its jamming. A packet starts in an “inactive” state by listening to the channel, and if a sufficient number of empty slots are heard, the packet becomes “active” in the sense of making transmission attempts. This “throttling” mechanism gives time for contention to decrease, thus allowing active packets to succeed. An active packet $u$ will make transmission attempts, but in contrast to updating based on feedback in each slot, $u$ monotonically reduces its sending probability as roughly $1/s_{u}$, where $s_{u}$ is the number of slots since $u$ became active. Packet $u$ also listens (when not transmitting), and if a sufficient number of empty slots are heard, it resets $s_{u}=1$. Here, they show expected constant throughput, with each packet making an expected number of transmission attempts that is polylogarithmic in $n+J$, where $J$ is the total number of jammed slots during which there exists at least one packet. Here, again, it is worth highlighting the related result by Chang et al. (ChangJP19, ) improves on this number of transmission attempts.
Recently, Bender et al. (bender:fully, ) showed how to achieve $\Omega(1)$ implicit throughput (recall Section 1.1.3) with $O({\texttt{poly}({\log n})})$ channel accesses (both sending and listening), despite an adaptive jamming adversary. This result applies in the channel setting with CD (i.e., ternary channel feedback), but with no control bits allowed, unlike the energy-efficient result of Bender et al. (BenderKPY18, ). The algorithm, named Low-Sensing Backoff, is based on multiplicative weights updates, and in that regard, it shares common ground with some prior results (such as (ChangJP19, ; richa:jamming2, ; richa:jamming3, ; richa:jamming4, ; awerbuch:jamming, ; ogierman2018sade, )). However, in comparison to these results, Low-Sensing Backoff greatly limits the frequency at which packets listen to the channel in order to update, while the lack of control bits greatly limits the amount of information they receive in each update. Analogous results are given for the cases where the number of packets is infinite, and where the adversary is reactive.
@@
Jurdziński et al. (DBLP:conf/waoa/JurdzinskiKL14, ) also consider a single sender and receiver, but in contrast to a parameterized adversary, they consider an unbounded adversary and a setting where packets of different sizes may arrive over time. If the adversary jams at any time during which a packet is transmitted, the transmission fails and the sender learns of this immediately. The authors derive results on the ratio between the throughput achieved by the algorithm and the throughput achieved by the best scheduling under identical arrival and error patterns; this is referred to as relative throughput. The authors give a deterministic online algorithm that achieves optimal relative throughput.
@@
An impossibility result is given by Bender et al. (bender2020contention, ), who show that with no-CD, no algorithm can guarantee a packet success in time linear in the system size with better than constant probability in the presence of jamming; this result holds even with a non-adaptive adversary. Subsequently, Chen et al. (DBLP:conf/podc/ChenJZ21, ) give an algorithm in the no-CD case that achieves a trade-off between the throughput and the amount of jamming tolerated. In the challenging case where a constant fraction of slots are jammed, $O(1/\log n)$ throughput is shown, which aligns with the above lower bound.
Summary.
@@
Perhaps unsurprisingly, randomization along with CD can provide constant throughput (in non-jammed slots) despite an unbounded adversary; even without CD, jamming can be tolerated, although throughput worsens slightly to $\Theta(1/\log n)$. When the adversary is bounded, a range of results exist. Under a saturated model, randomized results demonstrate constant throughput for a variety of network settings against a bounded adversary, so long as some weak information about the adversary’s power and $n$ is known. In contrast, deterministic results in the AQT setting are more nuanced, being parameterized by aspects of jamming rate and burstiness, and generally more pessimistic.
@@
Most results that tolerate jamming exhibit a sharp trade-off between the amount of listening devices perform and throughput achieved. On one hand, if each participant listens to the channel in every slot (except when sending), then constant throughput (in expectation or w.h.p. in non-jammed slots) is possible (bender:how, ; BenderFGY19, ; ChangJP19, ; awerbuch:jamming, ; richa:jamming2, ; richa:jamming3, ). Note that such results are not energy efficient, given that listening is roughly as costly as sending in practice (recall Section 1.1.3). On the other hand, if devices never listen to the channel, only sub-constant throughput is possible, and the best result is $O(1/\log n)$ throughput w.h.p. (DBLP:conf/podc/ChenJZ21, ). The recent result in (bender:fully, ) demonstrates that jamming-resistant contention resolution is possible with constant throughput and a polylogarithmic number of channel accesses (both sending and listening) without using control bits. However, it is unclear whether better energy-efficiency bounds are possible in the presence of jamming, perhaps under models with additional feedback permitted on the channel.